One side of a square lies along the straight line4š‘„ + 3š‘¦ = 26. The diagonalsĀ 

of the square intersect at the point (āˆ’2,3).

Find;

(a) The coordinates of the vertices of the square.

(b)The equation of the sides of the square which are perpendicular to the givenĀ 

line.


1
Expert's answer
2022-02-01T12:19:56-0500

a)

Among the coordinates of the equationĀ 4x +3y = 26Ā are:

(-4, 14), (-1, 10), ( 2, 6), (5, 2), (8, -2), (11, 6)



Considering the drawing, we can solve for the vertices of the square by finding the slope of one of the vertices against the center point and apply the same to the rest.

First, with reference to the given center point (2, 3) and following the conditions that (1)Ā one side of the square should be resting along the 4x+3y=26 lineĀ and (2)Ā the diagonals must be intersecting at the specified point (-2,3), we can see that points (5,2) & (-1, 10) are the only possible points that can be considered for such prerequisites as vertices.

The vertices are now represented by pointsĀ A, B, C, & D. And the center point asĀ V. Letā€™s get the slope of the line AV as our reference.

SlopeĀ m = (Y2 - Y1) / (X2 - X1)

where

(X2, Y2) =Ā (5,2)Ā & (X1, Y1) =Ā (-2, 3)

m = (2ā€“3) / (5-(-2))

m = -1 / 7

Since we already have the coordinates for vertices A & B, we can just apply the slope for points C & D.

Point C is located at the left side of the center point

X3Ā = (-2 - 7) =Ā -9

Y3Ā = (3-(-1)) =Ā 4

C (X3,Y3) = (-9, 4)

Point D is at the lower part of the center point, thus switching the slopeā€™s values .

X4Ā = (-2 - 1) =Ā -3

Y4Ā = (3 - 7) =Ā -4

D (X4,Y4) = (-3, -4)

The coordinates of the squareā€™s vertices are:Ā (5, 2), (-1, 10) (-9, 4) (-3, -4)


b)

Ā line 4š‘„ + 3š‘¦ = 26 "\implies y=-4x/3+26/3"

slope is -4/3

for perpendicular:

slope is 3/4


then:

for AD and point (5,2):

"2=3\cdot5/4+a \implies a=-1.75"

so,

"y=3x\/4-1.75"


for BC and point (-1,10):

"10=-3/4+a \implies a=10.75"

so,

"y=3x\/4+10.75"